OpenStudy (anonymous):

here is the question 5 and 1/2 - m = 2 and 5/6

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

here do this... 6 over 10 = 0.9 over n what is n??

OpenStudy (dumbcow):

convert mixed num to improper fractions \[\frac{11}{2} - m = \frac{17}{6}\] subtract 11/2 on both sides \[-m = \frac{17}{6} - \frac{11}{2}\] add fractions by finding common denominator of 6 \[-m = \frac{17}{6} - \frac{33}{6} = \frac{-16}{6}\] cancel out neg signs and reduce \[m = \frac{8}{3}\]

OpenStudy (anonymous):

wait is this for my problem...

OpenStudy (dumbcow):

yes the one you posted above those are the steps to solving for m